A platform is moving in upward direction with uniform velocity 40 m/s. When it reached at a height of 100 m from the ground a packet is dropped from it the total distance travelled by packet during its free fall is

Grade:12th pass

1 Answers

Mohammed Kasib Siddiqui
5 years ago
When the packet is dropped it acquires the velocity of the platform that is 40 m/s. As the platform is moving up therefore the packet also starts moving in the same direction as it had the velocity of the platform .
Now the maximum height reached by the particle so that its velocity becomes 0 when it had just been dropped will be-
v^2= u^2 - 2 gh
0= 1600- 2x10xh
h=80 m
Now, 
The total height acqiured by the particle will be the same that it will travel to ground...
Total height-
h = 100+80
h= 180m  ANS.